# Driving Referring Video Object Segmentation with Vision-Language Pre-trained Models

### v1

@@ III-B2 Language prompt-tuning
Language prompt-tuning aims to adapt the pre-trained language encoder for a comprehensive understanding of the referring expression. We append a set of language prompt tokens $\bm{P}_{e}\in\mathbb{R}^{m_{e}\times C_{e}}$ to the tokenized word embeddings. These tokens can learn the overall distribution of the referring expression data, and facilitate the language encoder modeling textual contexts to understand the referring expression.
@@
III-C Reference-guided visual encoding
Unlike many RVOS methods [7, 21] performing VL relation modeling only after feature extraction, we propose to inject the linguistic reference information into the visual encoder during feature extraction, serving as the first stage of VL relation modeling. As shown in Figure 4, we feed the language feature $\bm{F}_{e}$ into each vision transformer layer and calculate the cross-attention between $\bm{F}_{e}$ and the visual embeddings of each layer. To this end, we introduce a Multi-head Cross-Attention (MCA) operation in each vision transformer layer, which reuses the parameter of the existing MSA operation. Owing to the alignment nature between the visual and linguistic features, such a simple parameter-reusing MCA operation can effectively modulate the visual feature with the linguistic concept. The formulation of the attention operation in the $l$-th layer, i.e., Eqn. 3, is modified as follows:
@@ III-C Reference-guided visual encoding
III-D VL feature fusion
The VL Feature Fusion (VLFF) module, built on the deep VL feature $\mathcal{F}_{v}$ and $\bm{F}_{e}$, is used to associate high-level visual semantics with abstract linguistic concepts. As shown in Figure 6, it consists of a vision-to-language (V2L) propagation block, a language-to-vision (L2V) propagation block, and a fusion block. The V2L propagation block uses the global linguistic representation $\bm{x}_{e}$ as the query to calculate MCA with $\mathcal{F}_{v}$, enhancing the linguistic concepts in $\bm{x}_{e}$ relevant to the visual content. The L2V propagation block uses $\mathcal{F}_{v}$ as the query to calculate MCA with the word-level linguistic feature $\bm{F}_{e}$, enhancing the referred visual contents in $\mathcal{F}_{v}$. Herein skip connections are introduced for feature stability. Similar to [58], we rescale the decoder output in the V2L propagation block using a learnable factor $\bm{\alpha}\!\in\!\mathbb{R}^{C_{e}}$ with small initial values to preserve VL alignment.
With the enhanced VL features, VLFF calculates the pixel-wise cosine similarity between them, obtaining the similarity vectors $\mathcal{S}=\{\bm{s}^{t}\in\mathbb{R}^{N_{v}}\}_{t=T}^{T+T_{c}-1}$. After reshaping the enhanced visual features and similarity vectors, we concatenate them frame-by-frame for fusion. Finally, we use a projection layer to reduce the dimension of the fusion feature to $C$. We denote the fusion feature by $\mathcal{F}_{f}$.
@@ III-D VL feature fusion
III-E Spatial-temporal reasoning
The Spatial-Temporal Reasoning (STR) module aims to capture the dynamic vision contents (such as motions and appearance variations) related to the referring expression. As shown in Figure 4, it consists of the spatial-temporal encoder and the transformer decoder, which are used to model the spatial-temporal and vision-language relations, respectively, and repeated for $N$ times.
We devise a Cube-Frame Multi-head Self-Attention (CF-MSA) mechanism tailored for efficient and effective spatial-temporal encoding, as illustrated in Figure 7. Given the input features $\mathcal{F}_{f}\in\mathbb{R}^{T_{c}\times H\times W\times C}$, we first partition it into non-overlap cubes (i.e.,, 3D windows). For each token, we calculate its attention with itself and neighbor tokens belonging to the same frame as well as the same cube. Inspired by [59, 9], we shift the cube in space dimensions and recalculate attention within cubes again for cross-cube modeling. Compared with 3D SW/W-MSA proposed in [9] that models the spatial-temporal relations within 3D windows, CF-MSA further considers the intra-frame global spatial relation, which benefits for perceiving the target location more robustly and accurately. Compared with global MSA, CF-MSA omits the relation between two tokens which are across long distances in both space and time. Assuming the cube size is $T_{c}\!\times\!M_{w}\!\times\!M_{w}$, the computation complexity of the global MSA, 3D W-MSA, and our CF-MSA (without cube shift) operations [^1] Linear Projection and SoftMax are omitted in determining complexity. on $\mathcal{F}_{f}\in\mathbb{R}^{T_{c}\times H\times W\times C}$ are $\Omega(\text{MSA})=2(T_{c}HW)^{2}C$, Ω(3D $\Omega(\text{3D W-MSA})=2M_{w}^{2}T^{2}_{c}HWC$, and $\Omega(\text{CF-MSA})=2T_{c}HW((T_{c}-1)M_{w}^{2}+HW)C$, respectively.
The transformer decoder in STR models the relation between the shallow visual features and the linguistic feature, constituting the third stage of VL relation modeling. It introduces additional low-level semantic guidance from the shallow features, facilitating STR to understand the variations of the visual contents within a video clip.
@@ IV Experiments
TABLE I: Experimental results on Ref-YouTube-VOS and Ref-DAVIS17.
Algorithms Visual Backbone Ref-YouTube-VOS Ref-DAVIS17
Trained on Ref-Youtube-VOS alone
MTTR [6] Video-Swin-T 55.3 54.0 56.6
MANet [20] Video-Swin-T 55.6 54.8 56.5
ReferFormer [7] Video-Swin-T 56.0 54.8 57.3 55.8 51.8 59.8
SgMg [60] Video-Swin-T 58.9 57.7 60.0 56.7 53.3 60.0
SOC [61] Video-Swin-T 59.2 57.8 60.5 59.0 55.4 62.6
Ours (CLIP) ViT-B/16 59.7 57.9 61.5 60.3 56.7 64.0
Ours (CLIP) ViT-L/14 62.2 60.2 64.3 64.8 61.7 67.9
Ours (VLMo) VLMo-B/16 60.1 58.4 61.8 61.2 57.3 65.1
Ours (VLMo) VLMo-L/16 63.3 61.1 65.5 66.5 61.9 71.1
Pre-trained on Ref-COCO/+/g and fine-tuned on Ref-Youtube-VOS
R2VOS [3] Video-Swin-T 61.3 59.6 63.1
ReferFormer [7] Video-Swin-T 59.4 58.0 60.9 59.7 56.6 62.8
ReferFormer [7] Swin-L 62.4 60.8 64.0 60.5 57.6 63.4
ReferFormer [7] Video-Swin-B 62.9 61.3 64.6 61.1 58.1 64.1
VLT [62] Video-Swin-B 63.8 61.9 65.6 61.6 58.9 64.3
TCAD [31] Video-Swin-B 65.8 63.6 68.0 64.6 61.6 67.6
HTML [30] Swin-L 63.4 61.5 65.3 61.6 58.9 64.4
HTML [30] Video-Swin-B 63.4 61.5 65.2 62.1 59.2 65.1
OnlineRefer [27] Swin-L 63.5 61.6 65.5 64.8 61.6 67.7
OnlineRefer [27] Video-Swin-B 62.9 61.0 64.7 62.4 59.1 65.6
RefSAM [63] ViT-L 62.1 60.9 63.3 68.5 65.2 71.9
SgMg [60] Video-Swin-B 65.7 63.9 67.4 63.3 60.6 66.0
SOC [61] Video-Swin-B 66.0 64.1 67.9 64.2 61.0 67.4
Ours (CLIP) ViT-B/16 62.9 61.3 64.4 65.1 61.4 68.8
Ours (CLIP) ViT-L/14 66.0 63.6 68.3 68.2 64.6 71.8
Ours (VLMo) VLMo-B 63.1 61.5 64.7 65.5 60.7 69.4
Ours (VLMo) VLMo-L 67.6 65.3 69.8 70.2 66.3 74.1
IV-A Experimental settings
Datasets and metrics. We evaluate VLP-RVOS on Ref-Youtube-VOS [18], Ref-DAVIS17 [17], A2D-Sentences [16], and JHMDB-Sentences [16]. For Ref-Youtube-VOS and Ref-DAVIS17, region similarity $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average value $\mathcal{J}\&\mathcal{F}$ are used as metrics, following [18]. For A2D/JHMDB-Sentences, mAP, overall IoU, and mean IoU are used as metrics, following [16].
Implementation details. We test our algorithm using different VLP models, including ViT-B/16 CLIP, ViT-L/14 CLIP [11], VLMo-B, and VLMo-L [12]. $C$ is set to 256 to reduce computation complexity. $N$ is set to 4. We use the features from the $3^{rd}$, $7^{th}$, and $9^{th}$ layers for skip connections when ViT-B/16 CLIP or VLMo-B serves as the backbone which contains 12 visual transformer layers, and we use those from the $6^{th}$, $14^{th}$, and $18^{th}$ layers for skip connections when ViT-L/14 CLIP or VLMo-L serves as the backbone which contains 24 visual transformer layers.
During training, we freeze the VLP model and optimize the remaining parameters using AdamW [64] with a weight decay of $5\times 10^{-4}$ and a learning rate of $5\times 10^{-5}$. Specifically, for Ref-Youtube-VOS, we train the model on its training set alone and report results on its val set. We also try to pre-train our model on Ref-COCO/+/g [65, 66] and fine-tune it on Ref-Youtube-VOS, similar to [3, 23, 7]. For Ref-DAVIS17, we directly report the results of the models trained on Ref-Youtube-VOS, providing insights into cross-dataset generalization. For A2D/JHMDB-Sentences, we train our model on the A2D-Sentences training set alone following [16]. We use the Dice [67] and Focal [68] losses for end-to-end training, whose weights are tuned to be 5 and 2, respectively. For image training data [66, 65], we set $T_{c}$ to 1, similar to [7]. For video training data [16, 17, 18], we set $T_{c}$ to 6 and train our model with two consecutive clips sampled from the same videos for each iteration. Thus we can generate historical prompts from the former clip and feed them into the model when performing forward propagation on the latter clip. This training strategy enables our VLP-RVOS to learn to exploit the historical prior. During inference, $T_{c}$ is set to 6 by default to maintain consistency with the training settings. Besides, we directly output the predicted masks without post-processing. We will release our source codes.
TABLE II: Experimental results on A2D/JHMDB-Sentences. $\dagger$ denotes that the model is additionally pre-trained on Ref-COCO/+/g before being trained on A2D-Sentences.
Algorithms Visual Backbone A2D-Sentences JHMDB-Sentences
mAP (%) IoUOverall (%) IoUMean (%) mAP (%) IoUOverall (%) IoUMean (%)
ReferFormer [7] Video-Swin-B 55.0 78.6 70.3 43.7 73.0 71.8
SOC [61] Video-Swin-B 57.3 80.7 72.5 44.6 73.6 72.3
SgMg [60] Video-Swin-B 58.5 79.9 72.0 45.0 73.7 72.5
LBDT-4 [4] ResNet-50 47.2 70.4 62.1 41.1 64.5 65.8
TCAD [31] ResNet-50 0. 76.6 68.6 70.6 69.6
LoSh-R [69] Video-Swin-T 50.4 74.3 66.6 40.7 71.6 71.3
SOC [61] Video-Swin-T 50.4 74.7 66.9 39.7 70.7 70.1
ReferFormer [7] Video-Swin-S 53.9 77.7 69.8 42.4 72.8 71.5
OnlineRefer [27] Video-Swin-B 0. 79.6 70.5 73.5 71.9
Ours (CLIP) ViT-B/16 53.3 76.7 69.5 44.2 73.6 71.9
Ours (CLIP) ViT-L/14 59.4 84.0 75.3 46.0 77.9 75.9
Ours (VLMo) VLMo-B 53.9 78.5 72.7 44.6 73.7 72.3
Ours (VLMo) VLMo-L 63.1 86.2 77.7 47.1 78.3 76.6
TABLE III: FLOPs, speed, and $\mathcal{J}\&\mathcal{F}$. We measure FLOPs and speed over Ref-Youtube-VOS on RTX3090. Long videos are dropped to avoid GPU memory overflow. The models are pre-trained on Ref-COCO/+/g and fine-tuned on Ref-Youtube-VOS.
Algorithms Visual Inference FLOPs per frame Speed $\mathcal{J}\&\mathcal{F}$ (%)
Backbone Clip Length (G) (FPS) Ref-Youtube-VOS Ref-DAVIS17
ReferFormer [7] Video-Swin-T all video frames 72 59 59.4 59.6
ReferFormer [7] Video-Swin-B all video frames 132 35 62.9 61.1
ReferFormer [7] Swin-L all video frames 220 37 62.4 60.5
SgMg [60] Video-Swin-T all video frames 65 67 62.0 61.9
SgMg [60] Video-Swin-B all video frames 121 41 65.7 63.3
SOC [61] Video-Swin-T all video frames 43 64 62.4 63.5
SOC [61] Video-Swin-B all video frames 98 34 66.0 64.2
OnlineRefer [27] Swin-L 1 222 11 63.5 64.8
Ours (CLIP) ViT-B/16 6 (our default setting) 72 77 62.9 65.1
Ours (CLIP) ViT-L/14 6 (our default setting) 219 30 66.0 68.2
Ours (VLMo) VLMo-B 6 (our default setting) 61 55 63.1 65.5
Ours (VLMo) VLMo-L 6 (our default setting) 183 22 67.6 70.2
Ours (CLIP) ViT-B/16 all video frames 71 103 62.8 64.5
Ours (CLIP) ViT-L/14 all video frames 202 36 65.9 67.9
Ours (VLMo) VLMo-B all video frames 58 70 62.8 65.1
Ours (VLMo) VLMo-L all video frames 173 29 67.4 70.0
IV-B Comparison with state-of-the-art methods
Ref-Youtube-VOS & Ref-DAVIS17. Table I reports the experimental results on Ref-Youtube-VOS and Ref-DAVIS17. The models are divided into two groups based on the training protocol. Our model with VLMo-L achieves the best performance in all metrics in both groups on the two benchmarks. Compared with ReferFormer (Swin-L), HTML (Swin-L), OnlineRefer (Swin-L), and RefSAM (ViT-L), which also use the backbone pre-trained on images, our methods with ViT-L/14 or VLMo-L as the visual backbone improve the segmentation performance with substantial margins, demonstrating the effectiveness of our VLP-RVOS. Besides, our methods perform favorably against state-of-the-art methods using the backbone pre-trained on videos, which manifests the potential of our temporal modeling approach.
Besides, compared with other models, most of our models exhibit substantial advantages in cross-dataset evaluation, i.e., training on Ref-Youtube-VOS and testing on Ref-DAVIS17. For example, VLP-RVOS with VLMo-B and VLMo-L improves the performance by 1.3% and 6.0% in $\mathcal{J}\&\mathcal{F}$ on Ref-DAVIS17, respectively, compared with SOC which using Video-Swin-B as the backbone. The cross-dataset evaluation results highlight the strong generalization ability of our VLP-RVOS.
A2D-Sentences & JHMDB-Sentences. Table II presents the experimental results on the A2D-Sentences and JHMDB-Sentences datasets. Our methods using ViT-B/14 CLIP and VLMo-B perform favorably against recently proposed methods, including SOC [61], OnlineRefer [27], and ReferFormer [7]. Our algorithms using ViT-L/14 CLIP and VLMo-L outperform these models by large margins, even those additionally pre-trained on Ref-COCO/+/g.
Efficiency. Table III reports FLOPs, speeds, and $\mathcal{J}\&\mathcal{F}$. Our models with default setting ($T_{c}=6$) achieve real-time or near real-time speeds. Inference clip length $T_{c}$ also influences efficiency and performance. By setting $T_{c}$ to the video length like other algorithms, our VLP-RVOS can run much faster without obviously compromising segmentation performance. For example, VLP-RVOS with ViT-B/16 CLIP runs at 103 FPS with a $\mathcal{J}\&\mathcal{F}$ score of 62.8. Compared to other models with similar speeds, our models show comparable or superior performance on Ref-Youtube-VOS and Ref-DAVIS17, demonstrating the potential of applying VLP models to RVOS.
IV-C Ablation studies
In our ablation experiments, We use ViT-B/16 CLIP [11] as the VLP model and train all the variants on the Ref-Youtube-VOS training set alone.
TABLE IV: Ablation studies of each component on Ref-Youtube-VOS. LP, VP, TP, and HP denote the Language Prompt, Vision Prompt, Temporal Prompt, and Historical Prompt, respectively. Stage-1/2/3 denotes the three VL relation modeling stages. The best and second-best results are marked in bold and underline, respectively.
@@ IV-C Ablation studies
TABLE V: Comparisons of different adaptation and temporal modeling methods over our VLP-RVOS framework. Full fine-tuning means fine-tuning the entire vision encoder. Partial-$m$ means fine-tuning only the last $m$ layers of the vision encoder. PRTC denotes the Parameter-Reusing Temporal Capturing module. $\mathcal{J}\&\mathcal{F}$ is reported.
Adaptation Temporal modeling

### v2

@@ III-B2 Language prompt-tuning
Language prompt-tuning aims to adapt the pretrained language encoder to understand the referring expression. We append language prompt tokens $\bm{P}_{e}\!\in\!\mathbb{R}^{m_{e}\times C_{e}}$ to the tokenized word embeddings. These tokens learn the overall distribution of the referring expression data and facilitate the language encoder modeling textual contexts to understand the referring expression comprehensively.
@@
III-C Multi-stage VL relation modeling
Herein we present how to perform multi-stage VL relation modeling while and after feature extraction in VLP-RVOS.
III-C1 Reference-guided visual encoding during feature extraction
Unlike many RVOS methods [7, 22] performing VL relation modeling only after feature extraction, we propose to inject the linguistic reference information into the visual encoder during feature extraction, serving as the first stage of VL relation modeling. As shown in Figure 4, we feed the language feature $\bm{F}_{e}$ into each vision transformer layer and calculate the cross-attention between $\bm{F}_{e}$ and the visual embeddings of each layer. To this end, we introduce a Multi-head Cross-Attention (MCA) operation in each vision transformer layer, which reuses the parameter of the existing MSA operation. Owing to the alignment nature between the visual and linguistic features, such a simple parameter-reusing MCA operation can effectively modulate the visual feature with the linguistic concept. The formulation of the attention operation in the $l$-th layer, i.e., Eq. (3), is modified as follows:
@@ III-C1 Reference-guided visual encoding during feature extraction
III-C2 VL feature fusion after feature extraction
The VL Feature Fusion (VLFF) module, built on the deep VL feature $\mathcal{F}_{v}$ and $\bm{F}_{e}$, is used to associate high-level visual semantics with abstract linguistic concepts. As shown in Figure 6, it consists of a vision-to-language (V2L) propagation block, a language-to-vision (L2V) propagation block, and a fusion block. The V2L propagation block uses the global linguistic representation $\bm{x}_{e}$ as the query to calculate MCA with $\mathcal{F}_{v}$, enhancing the linguistic concepts in $\bm{x}_{e}$ relevant to the visual content. The L2V propagation block uses $\mathcal{F}_{v}$ as the query to calculate MCA with the word-level linguistic feature $\bm{F}_{e}$, enhancing the referred visual contents in $\mathcal{F}_{v}$. Herein skip connections are introduced for feature stability. Similar to [60], we rescale the decoder output in the V2L propagation block using a learnable factor $\bm{\alpha}\!\in\!\mathbb{R}^{C_{e}}$ with small initial values to preserve the alignment between VL features.
With the enhanced VL features, VLFF calculates the pixel-wise cosine similarity between them, obtaining the similarity vectors $\mathcal{S}=\{\bm{s}^{t}\in\mathbb{R}^{N_{v}}\}_{t=T}^{T+T_{c}-1}$. After reshaping the enhanced visual features and similarity vectors into 3D tensors, we concatenate them frame-by-frame for fusion. Finally, we use a projection layer to reduce the dimension of the fusion feature to $C$. We denote the fusion feature by $\mathcal{F}_{f}$.
@@ III-C2 VL feature fusion after feature extraction
III-C3 VL relation modeling with shallow features
We further model the VL relation between the shallow visual features and linguistic features in the STR module, which facilitates the STR module associating the changing visual contexts with linguistic concepts. It will be detailed in Section III-D.
III-D Spatial-temporal reasoning for RVOS
The Spatial-Temporal Reasoning (STR) module aims to capture the dynamic vision contents related to the referring expression, such as objects with varying appearances. As shown in Figure 4, it consists of the spatial-temporal encoder and the transformer decoder, which are used to model the spatial-temporal and vision-language relations, respectively, and repeated for $N$ times.
We devise a Cube-Frame Multi-head Self-Attention (CF-MSA) mechanism for efficient and effective spatial-temporal encoding, as illustrated in Figure 7. Given the input features $\mathcal{F}_{f}\in\mathbb{R}^{T_{c}\times H\times W\times C}$, we first partition it into non-overlap cubes (i.e., 3D windows). For each token, we calculate its attention with itself and neighbor tokens belonging to the same frame as well as the same cube. Inspired by [9], we shift the cube in space dimensions and recalculate attention within cubes again for cross-cube modeling. Compared with 3D SW/W-MSA proposed in [9] that models the spatial-temporal relations within 3D windows, CF-MSA further considers the intra-frame global spatial relation, which benefits for perceiving the target location more robustly and accurately. Compared with global MSA, CF-MSA omits the relation between two tokens across long spatial and temporal distances, facilitating model learning. Experimental results demonstrate that CF-MSA outperforms global MSA and 3D W-MSA for spatial-temporal encoding, leading to better RVOS performance.
Assuming the cube size is $T_{c}\!\times\!M_{w}\!\times\!M_{w}$, the computation complexity of the global MSA, 3D W-MSA, and our CF-MSA (w/o cube shift) operations [^1] Linear Projection and SoftMax are omitted in determining complexity. on $\mathcal{F}_{f}\!\in\!\mathbb{R}^{T_{c}\times H\times W\times C}$ are $\Omega(\text{MSA})=2(T_{c}HW)^{2}C$, Ω(3D $\Omega(\text{3D W-MSA})=2M_{w}^{2}T^{2}_{c}HWC$, and $\Omega(\text{CF-MSA})=2T_{c}HW((T_{c}-1)M_{w}^{2}+HW)C$, respectively. CF-MSA is comparable with 3D W-MSA but surpasses global MSA in efficiency.
The transformer decoder in STR models the relation between the shallow visual features and the linguistic features, constituting the third stage of VL relation modeling. It introduces additional low-level semantic guidance from the shallow visual features, facilitating STR to understand the variations of the visual contents within a video clip.
@@ IV Experiments
TABLE I: Ablation studies of each component on Ref-Youtube-VOS. LP, VP, TP, and HP denote the Language Prompt, Vision Prompt, Temporal Prompt, and Historical Prompt, respectively. Stage-1/2/3 denotes the three VL relation modeling stages.
@@
IV-A Experimental settings
Datasets and metrics. We evaluate VLP-RVOS on Ref-Youtube-VOS [18], Ref-DAVIS17 [17], A2D-Sentences [16], JHMDB-Sentences [16], and MeViS [19]. For Ref-Youtube-VOS, Ref-DAVIS17, and MeViS, region similarity $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average value $\mathcal{J}\&\mathcal{F}$ are used as metrics, following [18, 19]. For A2D/JHMDB-Sentences, mAP, overall IoU, and mean IoU are used as metrics, following [16].
Implementation details. We test our algorithm using different VLP models, including ViT-B/16 CLIP, ViT-L/14 CLIP [11], VLMo-B, and VLMo-L [12]. For ViT-B/16 CLIP, we enlarge the input image size from 224 to 352 and interpolate the pretrained positional embeddings. For ViT-L/14 CLIP, VLMo-B, and VLMo-L, we use the original input image sizes, which are 336, 384, and 384, respectively. $C$ is set to 256 to reduce computation complexity. $N$ is set to 4. During training, we freeze the VLP model and optimize the remaining parameters using AdamW [61] with a weight decay of $5\times 10^{-4}$ and a learning rate of $5\times 10^{-5}$. Specifically, for Ref-Youtube-VOS, we train the model on its training set alone and report results on its validation set. We also try to pretrain our model on Ref-COCO/+/g [62, 63] and fine-tune it on Ref-Youtube-VOS, similar to [7]. For Ref-DAVIS17, we directly report the results of the models trained on Ref-Youtube-VOS, providing insights into cross-dataset generalization. For A2D/JHMDB-Sentences, we train our model on the A2D-Sentences training set alone following [16]. For MeViS, we train the model on its training set alone, following [19]. We use the Dice [64] and Focal [65] losses for end-to-end learning, whose weights are tuned to be 5 and 2, respectively. For image training data [63, 62], we set $T_{c}$ to 1, similar to [7]. For video training data [16, 17, 18], we set $T_{c}$ to 6 and train our model with two consecutive clips sampled from the same videos for each iteration. Thus we can generate historical prompts from the former clip and feed them into the model when performing forward propagation on the latter clip, which allows our VLP-RVOS learning to exploit the historical prior. During inference, $T_{c}$ is set to 6 by default to maintain consistency with the training settings. We will release our source codes.
IV-B Ablation studies
We first conduct ablation studies to analyze our VLP-RVOS framework. We use ViT-B/16 CLIP [11] as the VLP model and train all the variants on Ref-Youtube-VOS alone for all ablation study experiments.
TABLE II: Comparisons of different adaptation and temporal modeling methods over our VLP-RVOS framework. Full fine-tuning means fine-tuning the entire vision encoder. Partial-$m$ means fine-tuning only the last $m$ layers of the vision encoder. PRTC denotes the Parameter-Reusing Temporal Capturing module. $\mathcal{J}\&\mathcal{F}$ is reported.
Tuning methods Temporal modeling methods